OpenStudy (anonymous):

At a school carnival, the diameter of the mat of a trampoline is 12 feet and the diameter of its metal frame is 14 feet. What is the length, in feet, of the metal frame that surrounds the trampoline? Use 3.14 and round your answer to the nearest tenth.

OpenStudy (rational):

draw a picture

OpenStudy (anonymous):

its like that

OpenStudy (rational):



OpenStudy (anonymous):

so whats the answer?

OpenStudy (rational):

radius = 7 use circumferencec formula to find the length of metal frame surrounding the trampoline

OpenStudy (rational):

you must be having the formula in ur notes look it up

OpenStudy (anonymous):

so I multiply with the 7

OpenStudy (rational):

yes, circumference = \(2\pi r = 2\times 3.14\times 7 = ?\)

OpenStudy (anonymous):

43.96

OpenStudy (rational):

\[\large \color{red}{\checkmark}\]